# Political Neutrality in AI is Impossible — But Here is How to Approximate it

### v1

@@ Abstract
AI systems often exhibit political bias, influencing users’ opinions and decision-making. While political neutrality—defined as the absence of bias—is often seen as an ideal solution for fairness and safety, this position paper argues that true political neutrality is neither feasible nor universally desirable due to its subjective nature and the biases inherent in AI training data, algorithms, and user interactions. However, inspired by Joseph Raz’s philosophical insight that “neutrality […] can be a matter of degree” (Raz, 1986), we argue that striving for some neutrality remains essential for promoting balanced AI interactions and mitigating user manipulation. Therefore, we use the term “approximation” of political neutrality to shift the focus from unattainable absolutes to achievable, practical proxies. We propose eight techniques for approximating neutrality across three levels of conceptualizing AI, examining their trade-offs and implementation strategies. In addition, we explore two concrete applications of these approximations to illustrate their practicality. Finally, we assess our framework on current large language models (LLMs) at the output level, providing a demonstration of how it can be evaluated. This work seeks to advance nuanced discussions of political neutrality in AI and promote the development of responsible, aligned language models.
@@ 1 Introduction
In recent years, large language models (LLMs) have been repeatedly shown to exhibit political bias (Feng et al., 2023; Röttger et al., 2024; Yang et al., 2024; Potter et al., 2024b). Moreover, recent studies have shown that interacting with politically biased LLMs can shape users’ political opinions and influence their decision-making (Fisher et al., 2024; Li, 2023; Hackenburg & Margetts, 2024; Durmus et al., 2024a; Potter et al., 2024b). Even so, these models are widely integrated in everyday applications, ranging from political news summarization (Zhang et al., 2024; Goyal et al., 2022) to detecting fake news (Chen & Shu, 2024), raising ethical concerns about independent opinion formation of users. A seemingly logical solution is to develop more politically neutral models (Rotaru et al., 2024; Lin et al., 2024; Durmus et al., 2024b; Pit et al., 2024b). However, in this paper we argue that true political neutrality is neither fully attainable nor universally desirable. This brings us to the critical question: If true political neutrality is unattainable, how should we address the problem of political bias in AI?
@@
In the context of this paper, political neutrality means being impartial, that is, not favoring some political viewpoints over others. The theoretical impossibility of achieving absolute political neutrality has been extensively explored in disciplines such as philosophy and political science (Merrill & Weinstock, 2014b; Iwasa, 2010; Raz, 1986). At the core of the challenge is the inherently subjective nature of political neutrality—what one cultural or ideological perspective perceives as neutral may be seen as biased by another (Perloff, 2018). Moreover, AI systems are fundamentally influenced by the biases embedded in their training data, algorithmic design, and deployment contexts (Yang & Roberts, 2021; Weidinger et al., 2022), making a technical achievement of political neutrality not easily achieved either.
Despite the theoretical and technical impossibility of achieving absolute political neutrality, we assert that approximations of political neutrality are both a practical and worthwhile endeavor. Inspired by Joseph Raz’s philosophical insight that “neutrality […] can be a matter of degree” and “one can deviate from complete neutrality to a greater or lesser extent”(Raz, 1986), we argue that striving for some neutrality remains essential for promoting balanced AI interactions and mitigating user manipulation. We use “approximation” to acknowledge the trade-offs inherent in each technique, recognizing that while they fall short of true neutrality, they bring us closer to it in varying degrees. This approach shifts the focus from an impossible ideal to a practical pursuit of different forms of impartiality.
To address this complexity, we introduce eight methods for approximating political neutrality across three levels conceptualizing AI—output, system, and ecosystem—illustrated in Figure 1, discussing possible methods for implementation and inherent trade-offs. Beyond proposing approximation techniques, we offer strategies to help system developers navigate the trade-offs involved in selecting appropriate methods based on specific application contexts (see example in Figure 2). Furthermore, we explore two practical applications of political neutrality, highlighting actionable steps toward impartiality. Finally, we provide empirical insights into the approximation techniques currently employed by LLMs at the output level, demonstrating how our framework can serve as a benchmark for future research. Our goal is to advance the NLP field by promoting more nuanced approaches to addressing political bias in LLMs and encouraging deeper exploration of effective political neutrality approximations in AI systems.
@@ 2 Political Neutrality in AI is Impossible
Defining Political Neutrality. Political philosopher John Rawls wrote that political neutrality may mean that “the state is not to do anything intended to favor or promote any particular comprehensive doctrine rather than another” (Rawls, 2005, p. 192). In the context of speech and the U.S., the most relevant legal text is the US Constitution’s First Amendment, which “prohibits the government from restricting speech based on the particular views expressed in that speech” (Congressional Research Service, 2024). While these definitions focus on the state, the abstract principles regarding the possibility and desirability of neutrality also applies to private actors such as AI developers.
@@
Political Neutrality Is Theoretically Impossible. Drawing on existing work in philosophy and political science, we argue that theoretically, political neutrality is impossible. First, we highlight the paradoxical nature of political neutrality. For example, for every political topic, it is impossible to avoid some kind of position-taking. In fact, the concept of political neutrality itself affirms specific concepts such as tolerance and civility (Rawls, 2005), and thus its basis prioritizes certain values and viewpoints over others. Further, there is no neutral point on the political spectrum—between left-leaning and right-leaning views lie moderate views, which are a political position themselves (see Iwasa (2010) for a related argument on the impossibility of policies that are equidistant to differing preferences). Even not taking any action or position, implicitly favors the stronger side, making achieving political neutrality through inaction impossible as well (Iwasa, 2010). This concept has been used to argue why neutrality in the form of inaction can exacerbate systemic issues, such as racism in the U.S. (Maye, 2022), or bias in international conflicts, where a neutral stance often benefits the stronger nation (Gavouneli, 2012).
@@
Lastly, evaluating political neutrality is often theoretically impossible as well. Consider, if the focus is on neutrality in the consequences of an action, this is difficult to evaluate due to the inherent uncertainty of outcomes. Alternatively, if we consider neutrality in the intent of an action, it is impossible to fully discern the true intent of a decision-maker (Merrill & Weinstock, 2014a). Therefore, from a philosophical standpoint, determining the ultimate success of political neutrality becomes infeasible.
@@
Political Neutrality in AI Is Technically Impossible. Beyond the theoretical infeasibility of political neutrality, there is the argument that achieving political neutrality in AI is currently technically impossible (LeCun, 2022; Potter et al., 2024a). This is primarily due to the process of creating AI models and reliance on human biased data and curation. For example, training datasets or those involved in RLHF may be biased—often unintentionally, but sometimes with the intention to shape the output—and thus induce bias in the model (Feng et al., 2023; Yang & Roberts, 2021). Lastly, the probabilistic nature of LLMs means that even if they were neutral in expectation, they could be biased in specific instances. Therefore, even though recent methods have reduced bias in AI along specific dimensions, completely removing bias remains an unsolved research challenge.
@@
Is Political Neutrality Desirable? Beyond the question of whether political neutrality is possible, another core question is whether it is desirable. There are moral, epistemological, and pragmatic reasons that make political neutrality desirable. Morally, political neutrality promotes individuals’ autonomy to make decisions, acknowledges that there are conflicting values, and equally respects all viewpoints (Merrill & Weinstock, 2014a). In terms of epistemological reasons, it is difficult to know which viewpoint is best, and people can reasonably disagree on viewpoints (Merrill & Weinstock, 2014a). Pragmatically, most LLMs are designed to provide information to a broader public, so political neutrality may be desirable, similar to some editorial standards in media (Wikipedia contributors, 2025), higher education (Kalven Committee, 1967), and government.
@@
However, there are also reasons why political neutrality may not be desirable, specially related to people’s preferences, companies’ free speech rights, effects on the information environment, and data quality. People may prefer models that express a political opinion. In fact, literature on motivated reasoning (Taber & Lodge, 2006) and confirmation bias (Nickerson, 1998) have shown that people prefer models that reinforce their own views (Sharma et al., 2023; Messer, 2025; Potter et al., 2024b). Related, attempts at political neutrality might be seen as censorship and reduce user agency. Further, private companies have free speech rights which encourages their additions to public discussion (Congressional Research Service, 2024). Additionally, political neutrality could negatively impact the information environment, potentially leading to information overload (Roetzel, 2019) if it presents all viewpoints, or suppressing free expression. Finally, data quality itself might differ by political viewpoints  (Potter et al., 2024a; Mosleh et al., 2024; Guess et al., 2019). Therefore, pursuing political neutrality may require incorporating lower-quality information (e.g., misinformation), which could compromise the reliability of a system.
@@
Given that true political neutrality is theoretically and technically impossible, we explore some methods of approximating political neutrality that could be practical and valuable depending on the context. These approximations involve methods that foster impartiality, bringing systems closer to neutrality. However, each technique varies in its proximity to true neutrality, offering developers the flexibility to select the most suitable approach for different contexts. By thoughtfully navigating trade-offs, AI developers can create systems that respect diverse viewpoints while promoting fairness, user autonomy, and trust.
@@ 3 Approximation of Political Neutrality in AI
We introduce eight approximation techniques across three levels: the output-level, which focuses on a model’s response; the system-level, which pertains to all input-output pairs of a single AI system; and the ecosystem-level, which spans all AI models in use. At each level, we define techniques to approximate political neutrality, discuss methods for implementation, and examine their inherent trade-offs. To compare these techniques, we use five key characteristics:
@@
  • Fairness: The technique promote impartial treatment of all viewpoints.
@@
For a discussion on why we selected these characteristics, see Section A.1. Table 1 compares these characteristics across approximation techniques, including formal mathematical definitions of each technique. For more details on these formal definitions see Section A.3.
@@
Table 1: Comparison of Approximations to Political Neutrality in AI Models. We define system $M$, input $x$, and output set $M(x)$. A user-preferred response is $y^{\star}$, with a semantic distance metric dist()()( ), threshold $k$, and biased output ${y_{i}}_{i=1}^{m}$ for $m$ reasonable viewpoints. Bias description for output is described as $b(i)$, and for systems as $B(i)$. Lastly, we define a system $M_{i}$ with bias $i$ and a user with bias $j$ as $U_{j}$, and two sets of metadata are $K$ and $L$. For more details, see Section A.3.
@@ 3.1 Output-Level
Approximation Technique: Refusal. Refusal involves deliberately refusing to respond to an input, a common practice in AI safety protocols (Han et al., 2024; Wen et al., 2024). Current refusal methods, designed to ensure safety, could be adapted to support political neutrality. These methods include fine-tuning on curated safety datasets (Wang et al., 2023), red-teaming to identify vulnerabilities (Hong et al., 2024), and reinforcement learning to optimize refusal decisions. System-level prompts, like those used by Anthropic (Anthropic, 2025c), offer another approach by instructing models to avoid subjective political questions. However, such prompts often struggle with nuanced cases involving implicit bias or coded language. A third option is detection systems that monitor inputs or outputs using static lists or dynamic classifiers. Examples include OpenAI’s Moderation API (OpenAI, ) and Meta’s Llama Guard (Inan et al., 2023), though they struggle with scoring political bias and setting refusal thresholds for political neutrality.
Tradeoffs. Refusal effectively avoids generating controversial or biased output, ensuring fairness and safety. Additionally, it is easy for users to understand that the model has refused to answer, leaving little room for misinterpretation. However, refusal often leads to user frustration, particularly when the model mistakenly applies it to safe inputs (Röttger et al., 2024). This tradeoff exemplifies the tension between providing helpful answers and avoiding potentially harmful or biased outputs. Refusal could also be harmful if a model does not provide certain information, e.g., preventing a minority group from knowing what their rights are. Lastly, only one-sided refusal of political responses could make a model biased at the system-level (Potter et al., 2024b).
@@
Tradeoffs. Avoidance can be safe and fair if the response is sufficiently distant from the direct answer. Also, it provides some information to the user, making it more useful than outright refusal. However, if the response is too disconnected from the user’s query, it risks frustrating or confusing the user and hindering their ability to obtain the desired answer. Moreover, avoidance can unintentionally introduce subtle biases. For instance, a factual response like “The current allocation for K-12 education is $30\%$” may be seen as endorsing the figure, despite merely stating a fact.
Approximation Technique: Reasonable Pluralism. Reasonable pluralism involves presenting all reasonable viewpoints in response to an input. This concept draws on Rawls’ work (1993), in which reasonable pluralism means that people in society hold diverse, yet reasonable and often conflicting, viewpoints. However, Rawls intentionally leaves the term “reasonable” vague, allowing for varied interpretations. Reasonable pluralism is also closely related to Overton pluralism proposed in Sorensen et al. (2024), when describing definitions of value pluralism. Related to value pluralism, RLHF methods have been shown to promote reasonable pluralism in models (Lake et al., 2024). Alternatively, aggregating outputs from diverse models with varying biases can achieve similar results (Feng et al., 2024). Pluralism can also be enhanced through targeted training or fine-tuning using feedback on missing perspectives from individuals across political views.
Tradeoffs. Reasonable pluralism offers the most comprehensive response by including all perspectives, ensuring fairness and providing users with a broad range of information. It also grants users full agency to access the information they seek and more. However, defining a “reasonable” viewpoint is contentious, and practical limitations prevent including all perspectives, introducing bias. Even presenting many sides, if not all, can lead to cognitive overload, as the responses tend to become quite verbose and could contain irrelevant information just to secure coverage. Lastly, presenting opposing perspectives equally can also lead to “both-sidesism,” where less credible viewpoints are treated as equally valid, potentially misleading users about their legitimacy (Aikin & Casey, 2022).
@@
One implementation approach, inspired by sociology, is “self-reflection” (Falk & Miller, 1998), where the model analyzes its own output to identify its biases. Techniques like chain-of-thought reasoning (Wei et al., 2022) or post-hoc rationalization (Madsen et al., 2021; Gurrapu et al., 2023) are methods that could assist a model in this task of analyzing the biases in its output. However, this self-analysis could inadvertently amplify existing biases instead of mitigating them. To address this, external systems could enhance transparency. While tools like the Gemini API (Google, ) and OpenAI’s Moderation API (OpenAI, ) assess safety, there is no classifier specifically designed to evaluate political bias, making it difficult to apply current safety methods to political neutrality effectively.
Tradeoffs. Output transparency gives users full agency by clearly labeling biases in the response, allowing them to assess and interpret the information themselves. It helps users understand the biases present, clarifying unsafe or partial content. However, the biased content itself can still pose risks, especially in sensitive contexts, as labeling bias does not eliminate its potential harm (Fisher et al., 2024).
Contextual Selection of Approximation Techniques. Given the tradeoffs of output-level approximation techniques, we propose two main approaches for selecting the appropriate technique for a given context: a static process and a dynamic process. A static process uses predefined principles to guide decisions, offering transparency and reproducibility. For example, a decision-tree (see Figure 2) guides the selection of an approximation based on user queries, this process provides interpretability but is also rigid and subject to design biases. For instance, the example in Figure 2 opts for providing partial information, avoidance, over outright refusal for inputs where it is infeasible to provide a balanced overview. While effective for straightforward inputs like “Where can I vote?”, nuanced queries (e.g., “Is climate change caused by human activity?”) pose challenges due to varying interpretations of the decision-questions and incomplete coverage of input diversity. In contrast, the dynamic process uses flexible mechanisms, such as aggregating diverse perspectives through democratic approaches (Ovadya et al., 2024a) such as RLHF, where users choose preferred responses. This method ensures inclusivity, alignment with user preferences, and personalization. While less interpretable, transparency can still be maintained through openly sharing aggregation methods. Dynamic processes are better suited for edge cases and adapting to diverse inputs but face challenges such as majority bias, scalability, and evolving social norms (Mill, 1859), making implementation resource-intensive.

### v2

@@ Abstract
AI systems often exhibit political bias, influencing users’ opinions and decisions. While political neutrality—defined as the absence of bias—is often seen as an ideal solution for fairness and safety, this position paper argues that true political neutrality is neither feasible nor universally desirable due to its subjective nature and the biases inherent in AI training data, algorithms, and user interactions. However, inspired by Joseph Raz’s philosophical insight that “neutrality […] can be a matter of degree” (Raz, 1986), we argue that striving for some neutrality remains essential for promoting balanced AI interactions and mitigating user manipulation. Therefore, we use the term “approximation” of political neutrality to shift the focus from unattainable absolutes to achievable, practical proxies. We propose eight techniques for approximating neutrality across three levels of conceptualizing AI, examining their trade-offs and implementation strategies. In addition, we explore two concrete applications of these approximations to illustrate their practicality. Finally, we assess our framework on current large language models (LLMs) at the output level, providing a demonstration of how it can be evaluated. This work seeks to advance nuanced discussions of political neutrality in AI and promote the development of responsible, aligned language models.
@@ 1 Introduction
In recent years, large language models (LLMs) have been repeatedly shown to exhibit political bias (Feng et al., 2023; Röttger et al., 2024a; Yang et al., 2024; Potter et al., 2024b). Moreover, recent studies have shown that interacting with politically biased LLMs can shape users’ political opinions and influence their decision-making (Fisher et al., 2025; Li, 2023; Hackenburg & Margetts, 2024; Durmus et al., 2024a; Potter et al., 2024b). Even so, these models are widely integrated in everyday applications, ranging from political news summarization (Zhang et al., 2024; Goyal et al., 2023) to detecting fake news (Chen & Shu, 2024), raising ethical concerns about independent opinion formation of users. A seemingly logical solution is to develop more politically neutral models (Rotaru et al., 2024; Lin et al., 2025; Durmus et al., 2024b; Pit et al., 2024). However, in this paper we argue that true political neutrality is neither fully attainable nor universally desirable. This brings us to the critical question: If true political neutrality is unattainable, how should we address the problem of political bias in AI?
@@
In the context of this paper, political neutrality means being impartial, that is, not favoring some political viewpoints over others. The theoretical impossibility of achieving absolute political neutrality has been extensively explored in disciplines such as philosophy and political science (Merrill & Weinstock, 2014; Iwasa, 2010; Raz, 1986). At the core of the challenge is the inherently subjective nature of political neutrality—what one cultural or ideological perspective perceives as neutral may be seen as biased by another (Perloff, 2018). Moreover, AI systems are fundamentally influenced by the biases embedded in their training data, algorithmic design, and deployment contexts (Yang & Roberts, 2021; Weidinger et al., 2022), making a technical achievement of political neutrality not easy achieved either.
Despite the theoretical and technical impossibility of achieving absolute political neutrality, we assert that approximations of political neutrality are both a practical and worthwhile endeavor. Inspired by Joseph Raz’s philosophical insight that “neutrality […] can be a matter of degree” and “one can deviate from complete neutrality to a greater or lesser extent”(Raz, 1986), we argue that striving for some neutrality remains essential for promoting balanced AI interactions and mitigating user manipulation. We use the term “approximation” to acknowledge the trade-offs inherent in each technique, recognizing that while they fall short of true neutrality, they bring us closer to it in varying degrees. This approach shifts the focus from an impossible ideal to a practical pursuit of different forms of neutrality.
We introduce eight methods for approximating political neutrality across three levels conceptualizing AI—output, system, and ecosystem—illustrated in Figure 1, discussing possible methods for implementation and inherent trade-offs. Beyond proposing approximation techniques, we offer strategies to help system developers navigate the trade-offs involved in selecting appropriate methods based on specific application contexts (see example in Figure 2). Furthermore, we explore two practical applications of political neutrality, highlighting actionable steps toward impartiality. Finally, we provide empirical insights into the approximation techniques currently employed by LLMs at the output level, demonstrating how our framework can serve as a benchmark for future research. Our goal is to advance the NLP field by promoting more nuanced approaches to addressing political bias in LLMs and encouraging deeper exploration of effective political neutrality approximations in AI systems.
@@ 2 Political Neutrality in AI is Impossible
Defining Political Neutrality. Political philosopher John Rawls wrote that political neutrality may mean that “the state is not to do anything intended to favor or promote any particular comprehensive doctrine [^1] Rawls describes a comprehensive doctrine as ones own views about life, right and wrong, or good and bad. (Rawls, 1993) rather than another” (Rawls, 1993, p. 192). In the context of speech and the U.S., the most relevant legal text is the US Constitution’s First Amendment, which “prohibits the government from restricting speech based on the particular views expressed in that speech” (Congressional Research Service, 2024). While these definitions focus on the state, the abstract principles regarding the possibility and desirability of neutrality also apply to private actors such as AI developers.
@@
Political Neutrality Is Theoretically Impossible. Drawing on existing work in philosophy and political science, we argue that theoretically, political neutrality is impossible. First, we highlight the paradoxical nature of political neutrality. For example, for every political topic, it is impossible to avoid some kind of position-taking. In fact, the concept of political neutrality itself affirms specific concepts such as tolerance and civility (Rawls, 1993), and thus its basis prioritizes certain values and viewpoints over others. Further, there is no neutral point on the political spectrum—between left-leaning and right-leaning views lie moderate views, which are a political position themselves (see Iwasa (2010) for a related argument on the impossibility of policies that are equidistant to differing preferences). Even not taking any action or position implicitly favors the stronger side, making achieving political neutrality through inaction impossible as well (Iwasa, 2010). This concept has been used to argue that neutrality in the form of inaction can exacerbate systemic issues, such as racism in the U.S. (Maye, 2022), or bias in international conflicts, where a neutral stance often benefits the stronger nation (Gavouneli, 2012).
@@
Lastly, evaluating political neutrality is often theoretically impossible as well. If the focus is on neutrality in the consequences of an action, this is difficult to evaluate due to the inherent uncertainty of outcomes. Alternatively, if we consider neutrality in the intent of an action, it is impossible to fully discern the true intent of a decision-maker (Merrill & Weinstock, 2014). Therefore, from a philosophical standpoint, determining the ultimate success of political neutrality becomes infeasible.
@@
Political Neutrality in AI Is Technically Impossible. Beyond the theoretical infeasibility of political neutrality, some argue that achieving political neutrality in AI is currently technically impossible (LeCun, 2022; Potter et al., 2024a). This is primarily due to the process of creating AI models and reliance on human biased data and curation. For example, training datasets or those involved in RLHF may be biased—often unintentionally, but sometimes with the intention to shape the output—and thus induce bias in the model (Feng et al., 2023; Yang & Roberts, 2021). Lastly, the probabilistic nature of LLMs means that even if they were neutral in expectation, they could be biased in specific instances. Therefore, even though recent methods have reduced bias in AI along specific dimensions, completely removing bias remains an unsolved research challenge.
@@
Is Political Neutrality Desirable? Beyond the question of whether political neutrality is possible, another core question is whether it is desirable. There are moral, epistemological, and pragmatic reasons that make political neutrality desirable. Morally, political neutrality promotes individuals’ autonomy to make decisions, acknowledges that there are conflicting values, and equally respects all viewpoints (Merrill & Weinstock, 2014). In terms of epistemological reasons, it is difficult to know which viewpoint is best, and people can reasonably disagree on viewpoints (Merrill & Weinstock, 2014). Pragmatically, political neutrality may be desirable for LLMs, just as it is in other domains that serve the public interest—such as media (Wikipedia contributors, 2025), higher education (Kalven Committee, 1967), and government.
@@
However, there are also reasons why political neutrality may not be desirable, specially related to people’s preferences, companies’ free speech rights, effects on the information environment, and data quality. People may prefer models that express a political opinion. In fact, people prefer models that reinforce their own views (Sharma et al., 2023; Messer, 2025; Potter et al., 2024b), in line with the literature on motivated reasoning (Taber & Lodge, 2006) and confirmation bias (Nickerson, 1998). Relatedly, attempts at political neutrality might be seen as censorship and reduce user agency. Further, private companies have free speech rights which, encourage their additions to public discussion (Congressional Research Service, 2024). Additionally, political neutrality could negatively impact the information environment, potentially leading to information overload (Roetzel, 2019) if it presents all viewpoints, or suppressing free expression. Finally, data quality itself might differ by political viewpoints  (Potter et al., 2024a; Mosleh et al., 2024; Guess et al., 2019). Therefore, pursuing political neutrality may require incorporating lower-quality information (e.g., misinformation), which could compromise the reliability of a system.
@@
Given that true political neutrality is theoretically and technically impossible, we explore some methods of approximating political neutrality that could be practical and valuable depending on the context. These approximations involve methods that promote aspects of neutrality. However, each technique varies in its proximity to true neutrality, offering developers the flexibility to select the most suitable approach for different contexts. By thoughtfully navigating trade-offs, AI developers can create systems that respect diverse viewpoints while promoting fairness, user autonomy, and trust.
@@ 3 Approximation of Political Neutrality in AI
We introduce eight approximation techniques across three levels: the output-level, which focuses on a model’s response; the system-level, which pertains to all input-output pairs of a single AI system; and the ecosystem-level, which spans all AI systems in use. At each level, we define techniques to approximate political neutrality, discuss methods for implementation, and examine their inherent trade-offs.
These techniques were chosen by examining related fields, like sociology, political science, and philosophy, which have long grappled with analogous questions around neutrality, bias, and representation. Drawing on insights from these disciplines allows us to ground our approximations in well-established debates and frameworks, even as we adapt them to the technical and practical constraints of AI systems.
To compare these techniques, we use five key characteristics:
@@
  • Fairness: The technique promotes impartial treatment of all viewpoints.
@@
For a discussion on why we selected these characteristics, see Section A.1. Table 1 compares these characteristics across approximation techniques, including formal mathematical definitions of each technique. For more details on these formal definitions see, Section A.3.
@@
Table 1: Comparison of Approximations of Political Neutrality in AI Models. We define system $M$, input $x$, and output set $M(x)$. A user-preferred response is $y^{\star}$, with a semantic distance metric dist()()( ), threshold $k$, and biased output $\{y_{i}\}_{i=1}^{m}$ for $m$ reasonable viewpoints. Bias description for output is described as $b(i)$, and for systems as $B(i)$. Lastly, we define a system $M_{i}$ with bias $i$ and a user with bias $j$ as $U_{j}$, and two sets of metadata are $K$ and $L$. For more details, see Section A.3.
@@ 3.1 Output-Level
Approximation Technique: Refusal. Refusal involves deliberately refusing to respond to an input, a common practice in AI safety protocols (Han et al., 2024; Wen et al., 2024). Current refusal methods, designed to ensure safety, could be adapted to support political neutrality. These methods include fine-tuning on curated safety datasets (Wang et al., 2024), red-teaming to identify vulnerabilities (Hong et al., 2024), and reinforcement learning to optimize refusal decisions. System-level prompts, like those used by Anthropic (Anthropic, 2025), offer another approach by instructing models to avoid subjective political questions. However, such prompts often struggle with nuanced cases involving implicit bias or coded language. A third option is detection systems that monitor inputs or outputs using static lists or dynamic classifiers. Examples include OpenAI’s Moderation API (OpenAI, 2025) and Meta’s Llama Guard (Inan et al., 2023), though they struggle with scoring political bias and setting refusal thresholds for political neutrality.
Tradeoffs. Refusal effectively avoids generating controversial or biased output, ensuring fairness and safety. Additionally, it is easy for users to understand that the model has refused to answer, leaving little room for misinterpretation. However, refusal often leads to user frustration, particularly when the model mistakenly applies it to safe inputs (Röttger et al., 2024b). This tradeoff exemplifies the tension between providing helpful answers and avoiding potentially harmful or biased outputs. Refusal could also be harmful if a model does not provide certain information, e.g., preventing a minority group from knowing what their rights are. Lastly, only one-sided refusal of political responses could make a model biased at the system-level (Potter et al., 2024b).
@@
Tradeoffs. Avoidance can be safe and fair if the response is sufficiently distant from the direct answer. Further, it provides some information to the user, making it more useful than outright refusal. However, if the response is too disconnected from the user’s query, it risks frustrating or confusing the user and hindering their ability to obtain the desired answer. Moreover, avoidance can unintentionally introduce subtle biases. For instance, a factual response like “The current allocation for K-12 education is $30\%$” may be seen as endorsing the figure, despite merely stating a fact.
Approximation Technique: Reasonable Pluralism. Reasonable pluralism involves presenting all reasonable viewpoints in response to an input. This concept draws on Rawls’ work (1993), in which reasonable pluralism means that people in society hold diverse, yet reasonable and often conflicting, viewpoints. Rawls contends that such diversity is an inherent feature of a liberal democratic society and, as such, must be accounted for in our political theories. However, Rawls intentionally leaves the term “reasonable” vague, allowing for varied interpretations. Reasonable pluralism is also closely related to Overton pluralism proposed in Sorensen et al. (2024), when describing definitions of value pluralism. Related to value pluralism, RLHF methods have been shown to promote reasonable pluralism in models (Lake et al., 2025). Alternatively, aggregating outputs from diverse models with varying biases can achieve similar results (Feng et al., 2024). Pluralism can also be enhanced through targeted training or fine-tuning from individuals with views across the political spectrum.
Tradeoffs. Reasonable pluralism offers the most comprehensive response by including many perspectives, ensuring fairness and providing users with a broad range of information. It also grants users full agency to access the information they seek and more. However, defining a “reasonable” viewpoint is contentious, and practical limitations prevent including all perspectives, introducing bias. Even presenting many sides, if not all, can lead to cognitive overload, as the responses tend to become quite verbose and could contain irrelevant information just to secure coverage. Lastly, presenting opposing perspectives equally can also lead to “both-sidesism,” where less credible viewpoints are treated as equally valid, potentially misleading users about their legitimacy (Aikin & Casey, 2022).
@@
One implementation approach, inspired by sociology, is “self-reflection” (Falk & Miller, 1998), where the model analyzes its own output to identify its biases. Techniques like chain-of-thought reasoning (Wei et al., 2022) or post-hoc rationalization (Madsen et al., 2021; Gurrapu et al., 2023) are methods that could assist a model in this task of analyzing the biases in its output. However, this self-analysis could inadvertently amplify existing biases instead of mitigating them. To address this, external systems could enhance transparency. While tools like the Gemini API (Google, 2024) and OpenAI’s Moderation API (OpenAI, 2025) assess safety, there is no classifier specifically designed to evaluate political bias, making it difficult to apply current safety methods to political neutrality effectively.
Tradeoffs. Output transparency gives users full agency by clearly labeling biases in the response, allowing them to assess and interpret the information themselves. It helps users understand the biases present, clarifying unsafe or partial content. However, the biased content itself can still pose risks, especially in sensitive contexts, as labeling bias does not eliminate its potential harm (Fisher et al., 2025).
Contextual Selection of Approximation Techniques. Given the tradeoffs of output-level approximation techniques, we propose two main approaches for selecting the appropriate technique for a given context: a static process and a dynamic process. A static process uses predefined principles to guide decisions, offering transparency and reproducibility. For example, a decision-tree (see an illustrative example in Figure 2) can guide the selection of an approximation technique based on user queries, which makes the process transparent, but is also rigid and subject to design biases. For instance, the example in Figure 2 opts for providing partial information or avoidance, over outright refusal for inputs where it is infeasible to provide a balanced overview. While effective for straightforward inputs like “Where can I vote?”, nuanced queries (e.g., “Is climate change caused by human activity?”) pose challenges due to varying interpretations of the decision-questions and incomplete coverage of input diversity. In contrast, the dynamic process uses flexible mechanisms, such as aggregating diverse perspectives through democratic approaches (Ovadya et al., 2024a) such as RLHF, where users choose preferred responses. This method ensures inclusivity, alignment with user preferences, and personalization. While less interpretable, transparency can still be maintained through openly sharing aggregation methods. Dynamic processes are better suited for edge cases and adapting to diverse inputs but face challenges such as majority bias, scalability, and evolving social norms (Mill, 1859), making implementation resource-intensive.